OpenStudy (anonymous):

Please help! One third of complement of an angle is five less than two times the angle. Find the angle. (Finally spelled it right) thanks!

OpenStudy (anonymous):

(round of applause for spelling) \[\frac{1}{3}(90-x) = 2x-5\] So, solving it: \[90-x= 6x-15\] \[7x=105\] \[x=15\]

OpenStudy (anonymous):

Thankss! I'm horrible at setting up equations..